WARNING: This web is not used anymore. Please use PDBService.MinuteS28August08 instead!

Minutes 28 August 08

  • Phone:
    • Gridka: Doris
    • Triumf: Andrew
    • RAL: Carmine
    • BNL: Carlos
    • NDGF: Olli
    • CNAF: Barbara

  • CERN:
    • 3D: Dirk, Maria, Luca, Lisa, Eva
    • ATLAS: Gancho

  • Today cannot join - reports sent by email:
    • PIC: databases running stable. CPU patch to be applied next week.

  • Sites status:
    • RAL: upgrade to + July CPU patch applied. 3rd node to be added to ATLAS RAC, foreseen by the end of the month. No storage upgrade is planned.
    • CNAF:
      • Upgrade to + July CPU patch scheduled for next week. This intervention will affect lhcb, atlas, fts and atlas lfc databases.
      • New storage will arrive mid september – some other databases will be moved to the new storage (no 3d databases).
      • Federico left CNAF. No plans to hire someone. Probably internal move to replace Federico.
    • Gridka: some problems after upgrading ATLAS and LHCb databases to – instances being rebooted. Cluster interconnect has been changed (Oracle support does not support direct cluster interconnect). CPU patch to be scheduled.
    • Triumf: FTS migrated to RAC environment. CPU patch applied to 3d and FTS databases. Storage upgrade affecting FTS intervention to be announced.
    • NDGF: CPU patch applied last Monday. Some load observed on the database from the clients.
    • BNL:
      • FTS to be patched (CPU patch) and downtime to upgrade storage firmware + update asmlib packages.
      • CPU patch applied on the 3d cluster.
      • Streaming among BNL and Slac: Atlas does not allow direct replication between cern and tier2 sites. Still being discussed.
      • Change default listener port – use both listeners until tnsnames.ora file is completely distributed.

  • CERN status:
    • July 08 CPU patch applied – rolling intervention on the downstream databases
    • Several interventions on the ATLAS Streams setup:
      • New schemas have been added and we have changed the filter rules used on the downstream database to support the new schemas.
      • Problems to add one of the new schemas in Taiwan. They are using a different character set and the use of transportable tablespaces is impossible. We have added a new rule on the apply side to skip new schema at Taiwan.
      • The use of TTS between different platforms (as old IN2P3 database is still use by Atlas) worked well.
      • Bug when dropping table fixed on ATLAS offline apply side by adding new rules to the apply processes.
      • PVSS setup is also completed.
    • LHCb Streams setup has been completed by adding the LHCb online database to the replication.
    • Lisa, new technical student – works with Eva on several Streams procedures automatizations and new features for the Streams monitoring

  • Upgrade to RedHat 5 (question by Triumf): not planned at any Tier1 site neither at CERN for the moment.

  • Experiment status:
    • ATLAS:
      • Planned to add 10 new configuration schemas to the replication between ATLAS online and offline databases.

  • Twiki documentation: PIC has requested to access to very specific CERN documentation in our twiki regarding database recovery. We need to understand why you want to access this specific information. Is there any additional documentation you would like to see in our twiki? Tier1 sites think that the documentation is ok.

  • Next meeting: 18th September

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2008-08-28 - EvaDafonte
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback